OpenStudy (anonymous):

find cos (x + y) if sin x = 5/13 and sin y = 4/5

OpenStudy (whpalmer4):

do you know the cosine sum of angles formula?

OpenStudy (anonymous):

I'm afraid I don't.

OpenStudy (whpalmer4):

cos(x + y) = cos x cos y - sin x sin y

OpenStudy (anonymous):

I don't know cosine of either of these though.

OpenStudy (whpalmer4):

ah, yes you do, you just don't realize it :-) Remember, sin^2 x + cos^2 x = 1 so if you know sin x, you can find cos x by rearranging that equation and substituting in the value you know for sin x

OpenStudy (anonymous):

cos^2 x = 1 - sin ^2 x

OpenStudy (whpalmer4):

if sin x = 5/13, then sin^2 x + cos^2 x = 1 (5/13)^2 + cos^2 x = 1 cos^2 x = 1-(5/13)^2 cos x = sqrt(1-(5/13)^2)

OpenStudy (whpalmer4):

and if you actually expand the stuff inside the square root, you'll get something for which the square root can be easily found

OpenStudy (whpalmer4):

I'm just too lazy to type the steps, sorry :-)

OpenStudy (whpalmer4):

oh, what the heck... sqrt of (1 - (25/169)) = sqrt (169/169 - 25/169) = sqrt (144/169) = 12/13

OpenStudy (anonymous):

12/13?

OpenStudy (anonymous):

good timing, haha

OpenStudy (whpalmer4):

yep! now do the same for the other one

OpenStudy (anonymous):

cos x^2 = sqrt 1 - sin x^2 = sqrt 1 - 4/5 = sqrt 25/25 - 16/25 = 9/25

OpenStudy (whpalmer4):

sqrt (9/25) = 3/5

OpenStudy (anonymous):

when you take the square root, should consider both solutions. I.e plus or minus?

OpenStudy (anonymous):

ah, didn't square root the last part. now how do i find cosine of x and y? do i just add 12/13 and 3/5?

OpenStudy (whpalmer4):

now you have all the pieces to use the identity I gave you

OpenStudy (whpalmer4):

cos (x+y) = cos x * cos y - sin x * sin y cos x = 12/13 cos y = 3/5 sin x = 5/13 sin y = 4/5

OpenStudy (whpalmer4):

@sourwing not necessary to consider the negative square root — these the side lengths of the triangle in the unit circle

OpenStudy (anonymous):

i got .246 when i plugged that into my calculator

OpenStudy (whpalmer4):

that's about right, the exact answer is 16/65

OpenStudy (whpalmer4):

should definitely do the exact answer in a problem like this

OpenStudy (anonymous):

alrighty thank you so much! i have two more with different identities that i don't know, if you wouldn't mind helping.

OpenStudy (anonymous):

@whpalmer4 can you help me on my question

OpenStudy (anonymous):

well, the sides length are positive, but their coordinate is not necessarily positive

OpenStudy (anonymous):

sin (x - y) if cos(x) = 8/17 and cos(y) = 3/5 tan(x - y) if csc(x) = 13/5 and cot(y) = 4/3

OpenStudy (whpalmer4):

okay, here are the identities: sin(x-y) = sin x cos y - cos x sin y tan x - y = (tan x - tan y)/(1 + tan x tan y) csc x = 1/sin x cot x = 1/tan x = cos x / sin x

OpenStudy (whpalmer4):

a handy listing of trig identities that I refer to often can be found at http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

ok, so if sin x ^2 + cos x^2 = 1, and i'm given cos, then sin x^2 = sqrt 1-cos(x)^2, right?

OpenStudy (whpalmer4):

close, sin x = sqrt (1 - cos^2 (x))

OpenStudy (anonymous):

square root 289/289 - 64/289 square root 225/289 15/17

OpenStudy (anonymous):

square root 1 - 9/25 square root 25/25 - 9/25 square root 16/25 4/5

OpenStudy (whpalmer4):

yes, those are both correct

OpenStudy (anonymous):

12/25 - 40/63?

OpenStudy (anonymous):

oops wait messed that one up

OpenStudy (anonymous):

12/25 - 120/289, can't figure out how to simplify 120/289

OpenStudy (anonymous):

i guess i cant lol since the only factors of 289 are 1, 17, and 289

OpenStudy (whpalmer4):

cos x = 8/17 cos y = 3/5 sin x = 15/17 sin y = 4/5 sin(x-y) = sin x cos y - cos x sin y = 15/17 * 3/5 - 8/17 * 4/5 = (45 - 32) / (5*17) = 13/85

OpenStudy (anonymous):

lol thank you

OpenStudy (whpalmer4):

how about that tan question? I'm fading quickly, act now before I fall asleep :-)

OpenStudy (aj14):

I have done everything and got to the last step and see the answer is 16/65 which is .246 but i get .2288 somehow.